# Seewo’s Submission to MLC-SLM: Lessons learned from Speech Reasoning Language Models

### v2

@@ 6 Conclusions
In this work, we present a comprehensive multi-stage training pipeline for multilingual conversational ASR, integrating LoRA, Chain-of-Thought data augmentation, and RLVR-based self-correction. Our approach achieves substantial improvements over the official challenge baselines. We summarize the results of our main version submitted to the evaluation set in table 6 and 7, to demonstrate the effectiveness of each component. Despite these advances, challenges remain in fully automating context-aware inference and further improving the granularity of self-correction. Future work will focus on enhancing inference efficiency, developing more sophisticated reward functions for reasoning, and extending our methods to broader real-world multilingual and conversational scenarios.

### v3

@@ 6 Conclusions
We presented a multi-stage training pipeline for multilingual conversational ASR, integrating LoRA, Chain-of-Thought data augmentation, and RLVR-based self-correction. Our approach achieves substantial improvements over the official challenge baselines. Table 6 and 7 summarize the results of our main submission to the evaluation set, demonstrating the effectiveness of each component. Despite these advances, challenges remain in fully automating context-aware inference and improving the granularity of self-correction. Future work will focus on enhancing inference efficiency, developing more sophisticated reward functions for reasoning, and extending our methods to broader real-world multilingual and conversational scenarios.